The water vapor formed after the complete combustion of 18cm3 of gaseous compound containing only C, H with O2, 162cm3 was counted. The volume of the remaining gas is 126cm3. If so, find the molecular formula of the compound. Explain with steps.
1
Expert's answer
2020-12-22T04:32:23-0500

CxHy + O2 → CO2 + H2O

CxHy + (x+y/4)O2 → xCO2 + (y/2)H2O

V(CxHy) = 18 cm3

V(H2O) = 162 cm3

V(CO2) = 126 cm3

"\frac{n(CO_2)}{n(C_xH_y )} = \frac{x}{1} = \frac{126}{18} \\

x = 7 \\

\frac{n(CO_2)}{n(H_2O)} = \frac{x}{\frac{y}{2}} \\

\frac{2x}{y} = \frac{126}{162} \\

 \frac{14}{y} = \frac{126}{162} \\

y = 18"

Answer: C7H18